OpenStudy (anonymous):

(0,6) 4x + 5y=2 please help what is the equation of the line y=

OpenStudy (shoba):

is the line perpendicular or parallel? 4x + 5y = 2 y = -4/5x + 1/2 y = mx + c from these equation m=-4/5 and c = 1/2 if it is parallel m=-4/5 if it is perpendicular m1m2 = -1 m2 = 5/4 then at this point (0,6)=(x1,y1) y - y1 = m(x-x1) - subtitude into this equation

OpenStudy (anonymous):

lol I am not sure

OpenStudy (shoba):

is the line perpendicular or parallel? 4x + 5y = 2 y = -4/5x + 2/5 y = mx + c from these equation m=-4/5 and c = 2/5 if it is parallel m=5/4 if it is perpendicular m1m2 = -1 m2 = 5/4 then at this point (0,6)=(x1,y1) y - y1 = m(x-x1) - substitute into this equation * these is the correct one ignore the previous

OpenStudy (anonymous):

My answer should be in a slope intercept form

OpenStudy (shoba):

then assume it to be parralel y - 6 = -4/5(x-0) y = -4/5x + 6 m=-4/5 which is the slope and c = 6 which is the intercept